# CADDreamer: CAD object Generation from Single-view Images

### v1

@@ 1 Introduction
To bridge the quality gap between shapes created by human designers and those generated by 3D generative models, this paper focuses on reconstructing Computer-Aided Design (CAD) models from sing-view images using diffusion-based generative models. CAD objects offer a higher level of abstraction in shape modeling and ensure structural integrity aligned with human perception. Consequently, generating CAD objects requires diffusion models to understand and interpret high-level structures and geometric primitives. However, single-view images capture only partial information about 3D shapes, necessitating the inference of complete CAD models comprising structured geometric primitives. Additionally, reconstructing the watertight boundary representation (B-rep) of CAD objects poses significant challenges due to inevitable noise and distortion in the generated geometric primitives. Even minor deviations in primitives can disrupt intersection edges, resulting in misaligned primitives or non-watertight B-reps that are unsuitable for manufacturing and product design.
@@
  • CADDreamer Framework: We propose CADDreamer, a two-stage framework for high-quality CAD reconstruction from single-view images. The first module leverages large-scale diffusion models [34], exhibiting superior generative capabilities and a powerful capacity to handle diverse geometric shapes. With cross-view and cross-domain attention mechanisms, the diffusion model ensures high consistency in both geometry and semantics across multiple views. However, the reconstructed mesh from multi-view images may inevitably suffer from distortions and noise. Even with semantic guidance, these inaccuracies can lead to failed computations of geometric primitive intersections, resulting in misaligned primitives or non-watertight B-reps. To address these issues, the second module of CADDreamer is designed to solve the associated geometric and topological challenges.
@@ 2 Related work
Single view reconstruction methods. Large-scale diffusion models [34], pre-trained on billion-scale data, have shown tremendous potential and robust performance in both 2D and 3D generation tasks. Numerous methods based on large-scale diffusion modules have been applied to tackle the problem of single-view 3D reconstruction and have achieved state-of-the-art performance [40, 38], such as Wonder3D [29], SyncDreamer [27], and LRM [10]. Despite their success, these approaches typically depend on low-level representations to reconstruct 3D shapes, which lack high-level shape comprehension. This often results in significant noise and distortions in the reconstructed meshes. In contrast, CADDreamer is designed to capture both low-level geometric representations and high-level structured representations of 3D shapes, including six types of primitives, to mitigate reconstruction noises and contribute to more compact geometric representations.
@@ 3 Method
As shown in Fig. 2, CADDreamer consists of two main modules: a multi-view generation module and a geometric and topological extraction module. First, the multi-view generation module takes a single-view RGB image to reconstruct a complete 3D triangular mesh and then decomposes this mesh into distinct patches, each corresponding to a primitive. Second, the geometric and topological extraction module refines the parameters of each primitive and performs CAD reconstruction through topology-preserving intersections between neighboring primitives. Sec. 3.1 details the multi-view generation module, while Sec. 3.2 elaborates on the geometric and topological extraction module.
@@ 3.1 Multi-view Generation Module
Thus, we developed a new multi-view generation module, as illustrated in Fig. 2. The entire process can be summarized in four steps. First, we use a popular normal predictor [2, 20] to generate a normal map from the single-view RGB image. Second, using the generated normal map as input, we fine-tune the popular cross-domain diffusion generative model, Wonder3D [29], to produce $m$ multi-view normal maps and $m$ multi-view semantic primitive maps, where $m=6$ views are generated in this paper. In these maps, the RGB values in the normal maps indicate normal directions, while those in each primitive map represent primitive types, including feature lines that separate neighboring primitives. To enhance generation performance, we fine-tune two separate VAE decoders for normal and primitive map generation, respectively. Third, we input multiple normal maps into NeuS [37] to generate a complete 3D mesh. Finally, we utilize feature lines to divide each primitive map into several 2D patches. These 2D patches are then back-projected onto the mesh, where a Graph Cut process is applied to merge the back-projected patches into a cohesive set of 3D mesh patches that represent different primitives.
@@
Graph Cut process. As shown in Fig. 3, each back-projected mesh patch is treated as an individual node, and connectivity edges are added between adjacent or overlapping patches, with the cosine similarity of the patch’s primitive parameters serving as the edge weight. Since the 3D patches derived from the same images correspond to different primitives, merging patches from the same images is considered an error in the graph cut methodology. Given that the number of primitives in CAD objects typically exceeds two, an initial cut threshold is set to divide the mesh into two surface patches. This threshold is then gradually increased to yield more connected components. The minimum cut threshold that results in the fewest errors is selected as the final cut threshold. Thus, through the Graph Cut process, the multi-view generation module effectively separates the constructed mesh into distinct surface patches, each representing a CAD primitive.
@@
Compared to previous models, our multi-view generation module incorporates two significant improvements. First, we introduce an off-the-shelf normal prediction model that converts the original images into normal maps, thereby alleviating the interference caused by image textures and shading variations. As a result, the impact of diverse textures and shading can be minimized, significantly reducing the problem’s complexity and contributing to a more robust reconstruction process. Second, we implement a new cross-domain generation strategy for multi-view normal and semantic primitive maps. By generating these maps together, we not only reconstruct a 3D mesh but also segment it into a structured representation of shapes, i.e., the five types of primitives illustrated in Tab. 1.
@@
$\vec{x},p$ : axis and center position
$\alpha$ : semi-angle
$h$ : height
@@
$r_{l}$ : major radius
$r_{s}$ : minor radius
@@
$p$ : center position
$r$ : radius
@@ 3.2 Geometric and Topological Extraction Module
  • First, we input the vertices of each mesh patch into the RANSAC [35] algorithm to extract the parameters for each primitive, as defined in Tab. 1.
  • Next, we introduce a novel geometric optimization algorithm to recover the topological relationships (e.g., intersections) and geometric relationships (e.g., parallelism, perpendicularity, and collinearity) between primitives, with detailed definitions provided in Tab. 1.
  • Third, we extract a topological representation from the segmented mesh patches, encompassing vertices, edges, and faces, as shown in Fig. 6(a).
@@
Primitive extraction. In the first step, we build on previous studies by using the RANSAC algorithm to estimate the parameters of primitives. Unlike direct least squares fitting, which can be significantly affected by outlier data, RANSAC helps to mitigate the impact of such outliers. However, due to the inevitable distortion in 3D mesh reconstruction from single-view images, the primitive parameters obtained from RANSAC may still be inaccurate. As a result, some essential topological and geometric relationships between the primitives might not be preserved. As illustrated in Fig. 4(a), the expected intersecting primitive surfaces may fail to intersect, leading to incorrect topological relationships and reconstruction failures for vertices and edges. Furthermore, since the parameters of primitives are based on points, lines, and directions, it is crucial to preserve key geometric relationships between these elements, such as collinearity among points and lines, as well as parallelism and perpendicularity among directions. As shown in Fig. 4(b), (c), and (d), neglecting these vital geometric relationships can result in completely erroneous intersection curves, which may cause gaps in the reconstructed CAD model. Therefore, in the next step, we aim to establish these important relationships between primitives before implementing the intersection reconstruction strategy.
@@
[…]
Figure 6: An illustration of the topology-preserving intersection strategy to achieve a watertight B-rep model with CAD vertices, edges, and faces. (a) depicts the topology derived from the mesh patches. (b) shows the CAD curves resulting from primitive intersections. (c) represents the reconstructed B-rep model through topology-preserving trimming.
@@
First, we identify all intersection relationships between the primitive surfaces. Second, for each intersection relationship, we select $k$ vertices on the mesh that are closest to the two intersecting primitive surfaces as stitching vertices. To minimize geometric calculation errors, these $k$ vertices are chosen from those located on patch boundaries, as shown in Fig. 5(b). Third, we project each stitching vertex onto the two primitive surfaces and minimize the distance between the corresponding projection points by optimizing the parameters of the primitive surfaces, as shown in Fig. 5(c-f).
@@
  • A collinear relationship implies that the position of primitive $A$ (denoted as $p_{A}$) can be expressed in terms of the parameters of primitive $B$: $p_{A}=p_{B}+\vec{x}_{B}*t$, where $\vec{x}_{B}$ is the axis direction of primitive $B$, as shown in Tab. 1. Therefore, to enforce collinearity constraints, we optimize the parameter $t$ instead of $p_{A}$.
@@
By employing the L-BFGS algorithm [6] to optimize the objective function, we effectively minimize the distances between primitive surfaces that should intersect and reconstruct the geometric relationships between primitives. Fig. 5 illustrates an example of four stitching points. As shown in Fig. 5 (d), by enforcing constraints to restore parallel and collinear relationships, the first step of the optimization process re-establishes the parallel relationship between the axes of the cylinders and planes. Subsequently, as the optimization progresses, we gradually restore the perpendicular and intersection relationships between the primitives, as shown in Fig. 5 (e-f).
Extraction of topological representation. In the third step, we extract a watertight topological representation from the mesh to guide the intersection of the primitives. Specifically, each 3D mesh patch corresponds to a topological face, while the edge curves adjacent to two patches are considered topological edges. Additionally, mesh vertices connected to more than two patches are regarded as topological vertices. The half-edge structure of the meshes provides directionality for the patch boundaries, and we inherit these directions as the orientations of the topological edges. This process enables us to extract a topological representation from the segmented mesh, including topological vertices, edges, and faces, as shown in Fig. 6(a). Since the reconstructed meshes are watertight, this topological representation is also watertight.
Figure 7: The segmentation results and reconstructed B-rep models. Electric blue ( ) represents segmentation boundaries and edges of B-rep objects. Other colors represents different primitives: plane ( ), cylinder ( ), cone ( ), sphere ( ), and torus ( ). Other colors represents different primitives: plane ( ), cylinder ( ), cone ( ), sphere ( ), torus ( ), and feature lines ( )
@@
First, we reconstruct the CAD curves for each face. Using Fig. 6(a) as an example, each topological edge’s neighborhood comprises two patches corresponding to two primitives. We then compute the intersection of these two primitives to obtain CAD curves. If multiple intersection curves exist between two primitives, we select the one with the closest projection distance to the corresponding topological edge. By applying this process to all topological edges, we generate the CAD curves for each one.
@@ 4 Experiments
Dataset. We began by training and evaluating CADDreamer on synthetic images, followed by testing it on real-world images. For the synthetic dataset, we curated 30,000 seamless CAD models from the ABC [15] and DeepCAD [39] datasets. These CAD objects were enhanced with textures and backgrounds, then rendered into multi-view images, resulting in a training set of 29,000 samples and a testing set of 1,000 samples. To maintain consistency, we adopted the camera configurations and image resolution settings used in Wonder3D [29]. Furthermore, we captured RGB images of real-world CAD objects using handheld devices to assess our model’s generalization performance on real-world objects.
Baselines. Currently, popular approaches for reconstructing B-reps fall into two categories: conventional sketch-extrude methods [4, 1, 43] and more recent intersection-based methods [28]. Because our dataset includes CAD models with a diverse range of primitives—such as spheres, tori, and cones—that cannot be reconstructed using conventional sketch-extrude methods [4, 1, 43], we focus on comparing intersection-based models, which can handle these varied primitive types. Our evaluation framework integrates state-of-the-art single-view reconstruction methods (e.g., SyncDreamer [27], LRM [10], CRM [38], and InstantMesh [40]) with B-rep reconstruction techniques like Point2CAD [28]. For benchmarking, we applied these single-view methods to synthetic images, generating meshes that were subsequently converted into B-reps using Point2CAD [28].
@@
Metrics. We evaluated generation quality across three aspects: meshes, primitives, and B-reps. First, an ideal reconstructed mesh should minimize reconstruction distortion and align closely with the original B-reps (i.e., primitives). To assess geometric alignment between the reconstructed mesh and the ground truth, we utilized two widely-used metrics: Chamfer Distance (CD) and Normal Consistency (NC) [8]. Second, incorrect segmentation results—particularly the omission of primitives—can cause significant fitting errors and reconstruction failures. To address this, we evaluated both the percentage of correctly classified mesh vertices (SEG(V)) and the accuracy of predicted primitive counts (SEG(P)). Additionally, fitting errors in primitives can disrupt correct intersection relationships, resulting in hanging faces or non-closed B-reps. To quantify these issues, we calculated the proportion of hanging faces (HF) in the resulting B-reps and measured the deviation of the reconstructed B-reps from the ground truth using Chamfer Distance (CD).
Settings. We used Wonder3D [29] as our pre-trained model to ensure more stable and faster convergence. All experiments were conducted on a single machine equipped with eight NVIDIA A100 GPUs (80GB each) and an AMD EPYC 7313 CPU. The details of the comparative analyses are presented below, while ablation studies and limitation analyses are provided in the supplementary materials.

### v2

@@ 1 Introduction
To bridge the quality gap between shapes created by human designers and those generated by 3D generative models, this paper focuses on reconstructing Computer-Aided Design (CAD) models from single-view images using diffusion-based generative models. CAD objects offer a higher level of abstraction in shape modeling and ensure structural integrity aligned with human perception. Consequently, generating CAD objects requires diffusion models to understand and interpret high-level structures and geometric primitives. However, single-view images capture only partial information about 3D shapes, necessitating the inference of complete CAD models comprising structured geometric primitives. Additionally, reconstructing the watertight boundary representation (B-rep) of CAD objects poses significant challenges due to inevitable noise and distortion in the generated geometric primitives. Even minor deviations in primitives can disrupt intersection edges, resulting in misaligned primitives or non-watertight B-reps that are unsuitable for manufacturing and product design.
@@
  • CADDreamer Framework: We propose CADDreamer, a two-module framework for high-quality CAD reconstruction from single-view images. The first module leverages large-scale diffusion models [35], exhibiting superior generative capabilities and a powerful capacity to handle diverse geometric shapes. With cross-view and cross-domain attention mechanisms, the diffusion model ensures high consistency in both geometry and semantics across multiple views. However, the reconstructed mesh from multi-view images may inevitably suffer from distortions and noise. Even with semantic guidance, these inaccuracies can lead to failed computations of geometric primitive intersections, resulting in misaligned primitives or non-watertight B-reps. To address these issues, the second module of CADDreamer is designed to solve the associated geometric and topological challenges.
@@ 2 Related work
Single-view Reconstruction Methods. Large-scale diffusion models [35], pre-trained on billion-scale data, have shown tremendous potential and robust performance in both 2D and 3D generation tasks. Numerous methods based on large-scale diffusion modules have been applied to tackle the problem of single-view 3D reconstruction and have achieved state-of-the-art performance [43, 40], such as Wonder3D [30], SyncDreamer [28], and LRM [11]. Despite their success, these approaches typically depend on low-level representations to reconstruct 3D shapes, which lack high-level shape comprehension. This often results in significant noise and distortions in the reconstructed meshes. In contrast, CADDreamer is designed to capture both low-level geometric representations and high-level structured representations of 3D shapes, including seven types of primitives, to mitigate reconstruction noises and contribute to more compact geometric representations.
@@ 3 Method
As shown in Figure 2, CADDreamer consists of two main modules: a multi-view generation module and a geometric and topological extraction module. First, the multi-view generation module takes a single-view RGB image to reconstruct a complete 3D triangular mesh and then decomposes this mesh into distinct patches, each corresponding to a primitive. Second, the geometric and topological extraction module refines the parameters of each primitive and performs CAD reconstruction through topology-preserving intersections between neighboring primitives. Sec. 3.1 details the multi-view generation module, while Sec. 3.2 elaborates on the geometric and topological extraction module.
@@ 3.1 Multi-view Generation Module
Thus, we develop a new multi-view generation module, as illustrated in Figure 2. The entire process can be summarized in four steps. First, we use a popular normal predictor [2, 21] to generate a normal map from the single-view RGB image. Second, using the generated normal map as input, we fine-tune the popular cross-domain diffusion generative model, Wonder3D [30], to produce $m$ multi-view normal maps and $m$ multi-view semantic primitive maps, where $m=6$ and feature lines have a width of three pixels in this paper. In these maps, the RGB values in the normal maps indicate normal directions, while those in each primitive map represent primitive types, including feature lines that separate neighboring primitives. To enhance generation performance, we fine-tune two separate VAE decoders for normal and primitive map generation, respectively. Third, we input multi-view normal maps into NeuS [39] for 3D mesh reconstruction. We remove the multi-view color inputs and associated texture reconstruction loss from NeuS, as CADDreamer does not require texture reconstruction. Finally, we utilize feature lines to divide each primitive map into several 2D patches. These 2D patches are then back-projected onto the mesh, where a Graph Cut process is applied to merge the back-projected patches into a cohesive set of 3D mesh patches that represent different primitives.
@@
Graph Cut Process. As shown in Figure 3, each back-projected mesh patch is treated as an individual node, and connectivity edges are added between adjacent or overlapping patches. Mesh patches sharing triangle boundaries are considered adjacent, while those with overlaps are overlapped. Using Efficient RANSAC [36], we extract primitive parameters and calculate their cosine similarity as edge weights. Since the 3D patches derived from the same images correspond to different primitives, merging patches from the same images is considered an error in the graph cut methodology. Given that the number of primitives in CAD objects typically exceeds two, an initial cut threshold is set to divide the mesh into two surface patches. This threshold is then gradually increased to yield more connected components. The minimum cut threshold that results in the fewest errors is selected as the final cut threshold. Thus, through the Graph Cut process, the multi-view generation module effectively separates the constructed mesh into distinct surface patches, each representing a CAD primitive.
@@
Compared to previous models, our multi-view generation module incorporates two significant improvements. First, we introduce an off-the-shelf normal prediction model that converts the original images into normal maps, thereby alleviating the interference caused by image textures and shading variations. As a result, the impact of diverse textures and shading can be minimized, significantly reducing the problem’s complexity and contributing to a more robust reconstruction process. Second, we implement a new cross-domain generation strategy for multi-view normal and semantic primitive maps. By generating these maps together, we not only reconstruct a 3D mesh but also segment it into a structured representation of shapes, i.e., the five types of primitives illustrated in Table 1.
@@
$\vec{x},p$ : axis and center position,
$\alpha$ : semi-angle, $h$ : height
@@
$r_{l}$ : major radius, $r_{s}$ : minor radius
@@
$p$ : center position, $r$ : radius
Relationships Determination
@@ 3.2 Geometric and Topological Extraction Module
  • First, we input the vertices of each mesh patch into the RANSAC [36] algorithm to extract the parameters for each primitive, as defined in Table 1.
  • Next, we introduce a novel geometric optimization algorithm to recover the topological relationships (e.g., intersections) and geometric relationships (e.g., parallelism, perpendicularity, and collinearity) between primitives, with detailed definitions provided in Table 1.
  • Third, we extract a topological representation from the segmented mesh patches, encompassing vertices, edges, and faces, as referred to Figure A1 in the supplement.
@@
Primitive Extraction. In the first step, we build on previous studies [36] by using the RANSAC algorithm to estimate the parameters of primitives. Unlike direct least squares fitting, which can be significantly affected by outlier data, RANSAC helps to mitigate the impact of such outliers. However, due to the inevitable distortion in 3D mesh reconstruction from single-view images, the primitive parameters obtained from RANSAC may still be inaccurate. As a result, some essential topological and geometric relationships between the primitives might not be preserved. As illustrated in Figure 4(a), the expected intersecting primitive surfaces may fail to intersect, leading to incorrect topological relationships and reconstruction failures for vertices and edges. Furthermore, since the parameters of primitives are based on points, lines, and directions, it is crucial to preserve key geometric relationships between these elements, such as collinearity among points and lines, as well as parallelism and perpendicularity among directions. As shown in Figure 4(b), (c), and (d), neglecting these vital geometric relationships can result in completely erroneous intersection curves, which may cause gaps in the reconstructed CAD model. Therefore, in the next step, we aim to establish these important relationships between primitives before implementing the intersection reconstruction strategy.
@@
First, we derive primitive intersection relationships from adjacent patches on segmented meshes: primitives intersect if and only if neighboring patches share boundaries. Second, for each intersection relationship, we sample $k$ boundary vertices closest to intersecting primitives as stitching vertices (Figure 5(b)). We uniformly sample candidate points along mesh segmentation boundaries (line-strings) and select $k$ points with minimal projection distances to the intersecting primitives as stitching vertices. Third, we project each stitching vertex onto the two primitive surfaces and minimize the distance between the corresponding projection points by optimizing the parameters of the primitive surfaces, as shown in Figure 5(c-f).
@@
  • A collinear relationship implies that the position of primitive $A$ (denoted as $p_{A}$) can be expressed in terms of the parameters of primitive $B$: $p_{A}=p_{B}+\vec{x}_{B}*t$, where $\vec{x}_{B}$ is the axis direction of primitive $B$, as shown in Table 1. Therefore, to enforce collinearity constraints, we optimize the parameter $t$ instead of $p_{A}$.
@@
In our implementation, the three geometric relationships are detected if the equations from Table 1 hold within a tolerance of 0.05. By employing the L-BFGS algorithm [7] to optimize the objective function, we effectively minimize the distances between primitive surfaces that should intersect and reconstruct the geometric relationships between primitives. Figure 5 illustrates an example of four stitching points. As shown in Figure 5 (d), by enforcing constraints to restore parallel and collinear relationships, the first step of the optimization process re-establishes the parallel relationship between the axes of the cylinders and planes. Subsequently, as the optimization progresses, we gradually restore the perpendicular and intersection relationships between the primitives, as shown in Figure 5 (e-f).
Extraction of Topological Representation. In the third step, we extract a watertight topological representation from the mesh to guide the intersection of the primitives. Specifically, each 3D mesh patch corresponds to a topological face, while the edge curves adjacent to two patches are considered topological edges. Additionally, mesh vertices connected to more than two patches are regarded as topological vertices. The half-edge structure of the meshes provides directionality for the patch boundaries, and we inherit these directions as the orientations of the topological edges. This process enables us to extract a topological representation from the segmented mesh, including topological vertices, edges, and faces (see Figure A1 in the supplementary material). Since the reconstructed meshes are watertight, this topological representation is also watertight.
Figure 6: The segmentation results and reconstructed B-rep models. Electric blue ( ) represents segmentation boundaries and edges of B-rep objects. Other colors represents different primitives: plane ( ), cylinder ( ), cone ( ), sphere ( ), and torus ( ).
@@
First, we reconstruct the CAD curves for each face. Referring to Figure A1(a) in the supplement, each topological edge’s neighborhood comprises two patches corresponding to two primitives. We then compute the intersection of these two primitives to obtain CAD curves. If multiple intersection curves exist between two primitives, we select the one with the closest projection distance to the corresponding topological edge. By applying this process to all topological edges, we generate the CAD curves for each one.
@@ 4 Experiments
Dataset. We begin by training and evaluating CADDreamer on synthetic images, followed by testing it on real-world images. For the synthetic dataset, we curate 30,000 seamless CAD models from the ABC [16] and DeepCAD [41] datasets. These CAD objects are enhanced with textures and backgrounds, then rendered into multi-view images, resulting in a training set of 29,000 samples and a testing set of 1,000 samples. To maintain consistency, we adopt the camera configurations and image resolution settings used in Wonder3D [30]. Furthermore, we capture RGB images of real CAD objects using handheld devices to assess our model’s generalization performance on real-world objects.
Baselines. Currently, popular approaches for reconstructing B-reps fall into two categories: conventional sketch-extrude methods [4, 1, 46] and more recent intersection-based methods [29]. Because our dataset includes CAD models with a diverse range of primitives—such as spheres, tori, and cones—that cannot be reconstructed using conventional sketch-extrude methods [4, 1, 46], we focus on comparing intersection-based models, which can handle these varied primitive types. Our evaluation framework integrates state-of-the-art single-view reconstruction methods (e.g., SyncDreamer [28], LRM [11], CRM [40], and InstantMesh [43]) with B-rep reconstruction techniques like Point2CAD [29]. For benchmarking, we apply these single-view generation methods to create meshes, and then use Point2CAD to segment the meshes into primitive shapes, estimate primitive parameters, and convert them into B-reps. Additional comparative assessments of the traditional sketch-extrude approach within point cloud inputs are documented in the supplementary material.
@@
Metrics. We evaluate generation quality across three aspects: meshes, primitives, and B-reps. First, an ideal reconstructed mesh should minimize reconstruction distortion and align closely with the original B-reps (i.e., primitives). To assess geometric alignment between the reconstructed mesh and the ground truth, we utilize two widely-used metrics: Chamfer Distance (CD) and Normal Consistency (NC) [9]. Second, incorrect segmentation results—particularly the omission of primitives—can cause significant fitting errors and reconstruction failures. To address this, we evaluate both the percentage of correctly classified mesh vertices (SEG(V)) and the accuracy of predicted primitive counts (SEG(P)). Additionally, fitting errors in primitives can disrupt correct intersection relationships, resulting in hanging faces or non-closed B-reps. To quantify these issues, we calculate the proportion of hanging faces (HF) in the resulting B-reps and measured the deviation of the reconstructed B-reps from the ground truth using Chamfer Distance (CD).
Settings. We use Wonder3D [30] as our pre-trained model to ensure more stable and faster convergence. All experiments are conducted on a single machine equipped with eight NVIDIA A100 GPUs (80GB each) and an AMD EPYC 7313 CPU. The details of the comparative analyses are presented below, while ablation studies and limitation analyses are provided in the supplementary materials.